OVHCloud Network Status

Current status
Legend
  • Operational
  • Degraded performance
  • Partial Outage
  • Major Outage
  • Under maintenance
FS#14792 — bhs2-20a/b-n56
Scheduled Maintenance Report for Network & Infrastructure
Completed
Both switches that are working in pair have rebooted due to an HA process dysfunction. We are working to have them up as fast as possible.

Update(s):

Date: 2015-09-24 13:25:58 UTC
All fex are up now, traffic has to go back on all customer interfaces.
Posted Sep 24, 2015 - 13:13 UTC